OpenStudy (anonymous):

Multiple choice!! x 2 3 4 5 P(X = x) 0.1 0.2 0.3 0.4 Find, E( Ix-4I ) A. 0 B. 0.6 C. 0.8 D. 1 E. 4 Answer is C. How come??

OpenStudy (zarkon):

\[\mathbb{E}[|x-4|]=\sum_{x=2}^5|x-4|P(X=x)\]